Please help me solve this problem (finding the pronumeral value) and explain the solution:
7(b-3)= -28

Answer by nyc_function(2741)

7(b-3)= -28
7b - 21 = -28
7b = -28 + 21
7b = -7
b = -7/7
b = -1